OpenStudy (anonymous):

What are the x-coordinates of the solutions to this system of equations? x2 + y2 = 16 y = x + 4 -4 and 0 4 and 0 -2 and -3 2 and -3

OpenStudy (imstuck):

Do you know how to use either the substitution method or the addition method to solve systems of equations?

OpenStudy (yanasidlinskiy):

I have a question. Is that ^ \[x^2\]

OpenStudy (anonymous):

solve \[x^2+\left( x+4 \right)^2=16~for~x\]

OpenStudy (anonymous):

yes @YanaSidlinskiy

OpenStudy (anonymous):

and i do know substitution method @IMStuck

OpenStudy (imstuck):

In the second equation you are told that y = x + y. You sub that y in into the first equation to eliminate the y and put the whole equation in terms of x, like this.\[x ^{2}+(x+4)^{2}=16\]Here's the substitution method!

OpenStudy (imstuck):

Actually you couldn't use the addition method here anyway because of the unlike x and y terms; one equation has them squared while the other does not. So any ideas on how to solve what I posted above?

OpenStudy (anonymous):

\[\left( a+b \right)^2=a^2+b^2+2*a*b\]